OpenStudy (anonymous):

A line contains points M(1, 3) and N(5, 0). What is the slope of MN?

OpenStudy (88):

(0-3)/(5-1)

OpenStudy (dayakar):

slope of a line passing through (x1,y1) = (1.3) and (x2,y2) =(5,0) is m= (y2-y1)/(x2-x1) can u do u it

OpenStudy (dayakar):

m= (0-3)/(5-1)

OpenStudy (88):

yes thats what i said

OpenStudy (anonymous):

i have to choose between A -4/3 B -3/4 C 3/4 D 4/3

OpenStudy (88):

a

OpenStudy (dayakar):

why don't u try

OpenStudy (dayakar):

B

OpenStudy (88):

i was about to say the same thing

OpenStudy (88):

i just had to look for a minute

OpenStudy (anonymous):

sorry im new with this and also cause i just started back to school havnt been in school for over a year so everything really blur to me .

OpenStudy (dayakar):

it's B

OpenStudy (88):

ok its ok because its been a blur to me to but i got back on my feet

OpenStudy (anonymous):

And i really have to answer all these questions fast cause if i dont it will auto submit my test everything i do is online .

OpenStudy (88):

u do k12

OpenStudy (anonymous):

its like that its called nccis 2020

OpenStudy (88):

oh

OpenStudy (anonymous):

yeah so can you guys help me with more

OpenStudy (88):

sure

OpenStudy (anonymous):

At work, Brett must check and record the internal temperature of the freezer on an hourly basis. When working properly, the temperature should remain constant over time. What word describes the slope of a line showing the temperature of the freezer as a function of time in hours when the freezer is working properly? The slope of the line will be.

OpenStudy (88):

The answer is a slope of ZERO

OpenStudy (anonymous):

thank you

OpenStudy (88):

np

OpenStudy (anonymous):

Line GH contains points G(–2, 6) and H(5, –3). What is the slope of ?

OpenStudy (88):

formula is RISE/RUN m= y2-y1/x2-x1

OpenStudy (anonymous):

Line GH contains points G(–2, 6) and H(5, –3). What is the slope of CH ? i ment

OpenStudy (88):

(-4,8)

OpenStudy (anonymous):

thats not a choice A -7/3 B -9/7 C -7/9 D -3/7

OpenStudy (88):

b

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The ice skating rink charges an hourly fee for skating and $3 to rent skates for the day. Gillian rented skates and skated for 3 hours and was charged $21. Which equation represents the cost, c(x), of ice skating as a function of x, the number of hours of skating?

OpenStudy (88):

np

OpenStudy (88):

c(3)=3×3+3=12

OpenStudy (anonymous):

A c(x) = 3x + 3 B c(x) = 6x + 3 C c(x) = 7x + 3 D c(x) = 8x + 3

OpenStudy (88):

c

OpenStudy (anonymous):

In order for the data in the table to represent a linear function with a rate of change of +5, what must be the value of m?

OpenStudy (88):

what are your choices

OpenStudy (anonymous):

m = 3 m = 8 m = 18 m = 33

OpenStudy (88):

m=18

OpenStudy (anonymous):

The slope of the line through the points is 2. Which statement describes how the slope relates to the height of the water in the pool?

OpenStudy (anonymous):

The height of the water increases 2 inches per minute. The height of the water decreases 2 inches per minute. The height of the water was 2 inches before any water was added. The height of the water will be 2 inches when the pool is filled.

OpenStudy (anonymous):

^ those are the choices

OpenStudy (88):

the height increases 2 inches per minute

OpenStudy (anonymous):

thank U

OpenStudy (88):

np

OpenStudy (anonymous):

Which equation represents the graphed function? y = –2x + 3 y = 2x + 3 y = x + 3 y = –x + 3

OpenStudy (88):

y=x+3

OpenStudy (anonymous):

my bad C. is y= 1/2x+3

OpenStudy (88):

y=2x+3

OpenStudy (anonymous):

okay , thanu u

OpenStudy (anonymous):

Which relationship has a zero slope?

OpenStudy (88):

i dont think i can help you with all of these because im home schooled and my middle school dance is today and i still have three more things to do

OpenStudy (anonymous):

okay its cool thank you so much though you really helped me alot